I need help with the following question. I tried the equation: .155*.324*e^(-4.49/.942) but I continue to get the wrong answer.Question 4, chap 131, sect 1. part 1 of 2 10 points A rectangular loop of area A is placed in a region where the magnetic field is perpendic- ular to the plane of the loop. The magnitude of the field is allowed to vary in time ac- cording to B = Boe., where Bo and t are constants. The field has a value of Bo when t is negative. Obtain a numerical value for E at t = 4.49 s when A = 0.155 1n2, Bo = 0.324 T. and T-0.912 s. Answer in units of mV Question 5, chap 131, sect 1. part 2 of 2 10 points For the values of A, Bo, and t given in part 1, what is the maximum value of et
Answer #1

1]We know that induced emf E = - d(BA)/dt = -A dB/dt = AB0e^(-t/T)/T = [0.155*0.324*e^(-4.49/0.942)]/0.942

= 0.0004537 V = 0.4537 mV answer

2] Emf E = AB0e^(-t/T)/T

maximum emf is also 0.4537 mV